# Spacetime symmetries and and geometric diffusion

### v1

@@ 1. Introduction
The Fokker–Planck (FP) equation describes diffusion processes, with applications in various domains of physics and engineering. In astrophysics, for instance, they are employed to model the theory of cosmic rays [lasuik2019]. Similarly, Fokker–Planck equations are utilized in plasma physics to analyze the effects of near-miss encounters between ions (heavy particles) and electrons [peeters2008].
@@
Building upon these ideas, A. Franchi and Y. Le Jan [franchi_jan] utilize the bundle $\mathcal{SO}^{+}(M)$ of direct pseudo-orthonormal frames, with fibers modelled on the special Lorentz group (and having their first element in the positive half of the unit pseudo-sphere in the tangent space) to extend the concept of relativistic diffusion to general Lorentzian manifolds. This approach defines a Stratonovich stochastic differential equation that takes values in the $\mathcal{SO}^{+}(M)$ group, similar to a Langevin equation. When projected, this equation naturally generates a diffusion process on the mass shell. Moreover, by following what is done in the Riemannian case, we can induce a pure diffusion on $M$ through a pullback operation, which results in a Fokker–Planck-type equation.
@@
In this paper, we show that the perspective provided by the frame bundle $\mathcal{SO}^{+}(M)$ helps to incorporate spacetime symmetries and define symmetric solutions to the Fokker–Planck equations both in the frame and observer bundle description. The careful way of incorporating symmetries into the equations has been exemplified in the case of the Vlasov equation—as demonstrated in works such as [sarbach2013relativistic] by Sarbach and Zannias—particularly on the mass shell. Here, we present a generalization that encompasses the previous result as a special case while also enabling the incorporation of diffusion.
@@
The structure of the paper is as follows: After this introduction, in Section 2, we present the necessary geometric structures necessary to describe diffusion processes in the frame bundle of a spacetime. We discuss the infinitesimal generator $\mathcal{L}$ of the Franchi–Le Jan-process, and we prove that, when the isometry group of the spacetime is non-trivial, its symmetry-reduced version is consistent. We also elucidate the connection of $\mathcal{L}$ with the corresponding FP equation (symmetric or not) on the $UM$ bundle. In Section 3, following a reasoning similar to that of Elhers in [ehlers, ehlers1973survey], we present a justification of the FP equations based on the conservation of the average number of particle world lines crossing any Cauchy hypersurface. We also define the particle current density, entropy current, and the energy-momentum tensor, along with some of their most important properties. In Section 4, we provide the explicit equations in several relevant spacetimes: flat Friedmann–Lemaître–Robertson–Walker cosmologies, the exterior Schwarzschild spacetime, and the Nariai spacetime. Finally, in Section 5, we present our conclusions and outline some future lines of research.
@@ 2.3. Symmetries
In addition to these conservation laws, as discussed in [KN, mok1979complete], given any vector field $\xi\in\mathfrak{X}(M)$, there exists a complete lift (natural lift) of $\xi$ to $\mathcal{F}(M)$ which we will denote as $\xi^{\mathrm{c}}\in\mathfrak{X}(\mathcal{F}(M))$. The complete lift is characterised by the following three properties:
@@ 4. Some relevant spacetimes
In this section, we use the results obtained in the previous sections to obtain the Fokker–Planck equations for one-particle distribution functions preserving the underlying symmetries in two physically relevant spacetimes.
@@ 4.2. Exterior Schwarzschild spacetime
$M=\mathbb{R}\times(r_{s},\infty)\times\mathbb{S}^{2}$
, and, in spherical coordinates, the line-element is given by
@@ 5. Conclusions and comments
In this paper, we have extensively discussed the relationship between the symmetry properties of a spacetime $(M,g)$ and relativistic diffusion processes on two relevant bundles over $M$: the (restricted) orthonormal frame bundle $\mathcal{SO}^{+}(M)$ and the observer bundle $U^{m}M$ of arbitrary positive mass $m$. We have demonstrated that the complete lift of a Killing vector field of $(M,g)$ commutes with both fundamental and basic vector fields within $\mathcal{SO}^{+}(M)$. This observation indicates that the symmetries inherent to the spacetime $(M,g)$ also preserve the generator of the diffusion processes on both $\mathcal{SO}^{+}(M)$ and $U^{m}M$. Furthermore, we have identified the conditions that satisfy symmetric solutions to the Fokker–Planck diffusion equation on $\mathcal{SO}^{+}(M)$ and $U^{m}M$, and we have detailed how to establish the connection between them, thereby extending previous findings in the literature.
To establish connections with numerous physical applications, we have demonstrated that within the framework provided by $U^{m}M$, the dynamics of the particle system can be characterized by the one-particle distribution function $f$, which obeys the Fokker–Planck equation ensuring the conservation of the average of particle trajectories passing through a given Cauchy surface. This conservation principle can also be derived through the vanishing divergence of the current density vector field $\mathsf{J}$ associated with $f$. Additionally, we have proven that the divergence of the entropy current $\mathsf{S}$ is non-negative, and that the divergence of the energy-momentum tensor field $\mathsf{T}$ is proportional to the particle current density, consistent with the expected behavior of a diffusion process.
While all the results presented in this work can be established without resorting to symplectic techniques, we have chosen to follow this approach anticipating future uses. For instance, it streamlines the incorporation of additional physical information, as is the case of electromagnetic fields: by simply modifying the symplectic form $\mathrm{d}\alpha$ to $\mathrm{d}\alpha+qF$, where $q$ represents the particle charge and $F$ denotes the Faraday tensor describing the electromagnetic field in spacetime, we can seamlessly integrate electromagnetic effects into our analysis.
Finally, we have illustrated the results using several physically relevant spacetimes, spanning both cosmological and astrophysical applications.

### v2

# Spacetime symmetries and geometric diffusion

@@ 1. Introduction
The Fokker–Planck (FP) equation describes diffusion processes, with applications in various domains of physics and engineering. In astrophysics, for instance, it is employed to model the theory of cosmic rays [lasuik2019]. Similarly, Fokker–Planck equations are utilized in plasma physics to analyze the effects of near-miss encounters between ions (heavy particles) and electrons [peeters2008].
@@
Building upon these ideas, A. Franchi and Y. Le Jan [franchi_jan] utilized the bundle $\mathcal{SO}^{+}(M)$ of direct pseudo-orthonormal frames, with fibers modelled on the special Lorentz group (and having their first element in the positive half of the unit pseudo-sphere in the tangent space) to extend the concept of relativistic diffusion to general Lorentzian manifolds. This approach defines a Stratonovich stochastic differential equation that takes values in the $\mathcal{SO}^{+}(M)$ group, similar to a Langevin equation. When projected, this equation naturally generates a diffusion process on the mass shell. Moreover, by following what is done in the Riemannian case, we can induce a pure diffusion on $M$ through a pullback operation, which results in a Fokker–Planck-type equation.
Recent works related to these methodologies include Serva’s study [serva2021brownian], which diverges from the approaches of Franchi and Le Jan by focusing on massless particles and constructing Lorentz invariant processes. Additionally, the work by Andra and Rosyid [AndraDoniRosyid] investigates relativistic diffusion incorporating both diffusion and friction within the framework of $f(R)$-gravity. Complementing these, the paper by Haba [haba2017thermodynamics], also within $f(R)$-gravity theory, examines the cosmological implications of relativistic diffusion, providing a comprehensive framework for understanding its impact on the universe’s large-scale structure and thermodynamic properties.
@@
In this paper, we show that the perspective provided by the frame bundle $\mathcal{SO}^{+}(M)$ helps to incorporate spacetime symmetries and define symmetric solutions to the Fokker–Planck equations both in the frame and observer bundle description. The careful way of incorporating symmetries into the equations has been exemplified in the case of the Vlasov [^1] In this context the Vlasov (Liouville) equation describes the evolution of the one-particle distribution function within relativistic kinetic theory, asserting that particles are conserved in phase space. Fokker-Planck reduces to Vlasov equation when the strength of the diffusion process $\sigma$ vanishes. (Liouville) equation—as demonstrated in works such as [sarbach2013relativistic] by Sarbach and Zannias—particularly on the mass shell. Here, we present a generalization that encompasses the previous result as a special case while also enabling the incorporation of diffusion.
@@
The structure of the paper is as follows: After this introduction, in Section 2, we present the geometric structures necessary to describe diffusion processes in the frame bundle of a spacetime. We discuss the infinitesimal generator $\mathcal{L}$ of the Franchi–Le Jan-process, and we prove that, when the isometry group of the spacetime is non-trivial, its symmetry-reduced version is consistent. We also elucidate the connection of $\mathcal{L}$ with the corresponding FP equation (symmetric or not) on the $UM$ bundle. In Section 3, following a reasoning similar to that of Elhers in [ehlers, ehlers1973survey], we present a justification of the FP equations based on the conservation of the average number of particle world lines crossing any Cauchy hypersurface. We also define the particle current density, entropy current, and the energy-momentum tensor, along with some of their most important properties. In Section 4, we provide the explicit equations in several relevant spacetimes: flat Friedmann–Lemaître–Robertson–Walker cosmologies, the exterior Schwarzschild spacetime, and the Nariai spacetime. Finally, in Section 5, we present our conclusions and outline some future lines of research.
@@ 2.3. Symmetries
In addition to these conservation laws, as discussed in [KN, mok1979complete], given any vector field $\xi\in\mathfrak{X}(M)$, there exists a complete lift (natural lift) of $\xi$ to $\mathcal{F}(M)$ which we will denote as $\xi^{\mathrm{c}}\in\mathfrak{X}(\mathcal{F}(M))$. The complete lift is characterized by the following three properties:
@@ 4. Some relevant spacetimes
In this section, we use the results presented in the previous sections to obtain the Fokker–Planck equations for one-particle distribution functions preserving the underlying symmetries in two physically relevant spacetimes.
@@ 4.2. Exterior Schwarzschild spacetime
$M=\mathbb{R}\times(r_{s},\infty)\times\mathbb{S}^{2}\,,$
and, in spherical coordinates, the line-element is given by
@@ 5. Conclusions and comments
In this paper, we have extensively discussed the relationship between the symmetry properties of a spacetime $(M,g)$ and relativistic diffusion processes on two relevant bundles over $M$: the (restricted) orthonormal frame bundle $\mathcal{SO}^{+}(M)$ and the observer bundle $U^{m}M$ of arbitrary positive mass $m$. We have demonstrated that the complete lift of a Killing vector field of $(M,g)$ commutes with both the fundamental and basic vector fields within $\mathcal{SO}^{+}(M)$. This observation indicates that the symmetries inherent to the spacetime $(M,g)$ also preserve the generator of the diffusion processes on both $\mathcal{SO}^{+}(M)$ and $U^{m}M$. Furthermore, we have identified the conditions that the symmetric solutions to the Fokker–Planck diffusion equation satisfy on $\mathcal{SO}^{+}(M)$ and $U^{m}M$, and we have detailed how to establish the connection between them, thereby extending previous findings in the literature.
To connect our results with physical applications, we have demonstrated that, within the framework provided by $U^{m}M$, the dynamics of the particle system can be characterized by the one-particle distribution function $f$, which obeys the Fokker–Planck equation ensuring the conservation of the average of particle trajectories passing through a given Cauchy surface. This conservation principle can also be derived through the vanishing divergence of the current density vector field $\mathsf{J}$ associated with $f$. Additionally, we have proven that the divergence of the entropy current $\mathsf{S}$ is non-negative, and that the divergence of the energy-momentum tensor field $\mathsf{T}$ is proportional to the particle current density, consistent with the expected behavior of a diffusion process.
Finally, we have illustrated the results by using several physically relevant spacetimes, spanning both cosmological and astrophysical applications.
While all the results presented in this work can be established without resorting to symplectic techniques, we have chosen to follow this approach anticipating future uses. For instance, it streamlines the incorporation of additional physical information, as is the case of electromagnetic fields: by simply modifying the symplectic form $\mathrm{d}\alpha$ to $\mathrm{d}\alpha+qF$, where $q$ represents the particle charge and $F$ denotes the Faraday tensor describing the electromagnetic field in spacetime, we can seamlessly integrate electromagnetic effects into our analysis. It may also be interesting to analyze recent generalizations of the Sasaki metric proposed in [kapsabelis2024finsler] within the context of modified gravity.